# Three Dimensional Topological Field Theories and Nahm Sum Formulas

### v1

@@ Abstract
It is known that a large class of characters of 2d conformal field theories (CFTs) can be written in the form of a Nahm sum. In Zagier:2007knq , D. Zagier identified a list of Nahm sum expressions that are modular functions under a congruence subgroup of $SL(2,\mathbb{Z})$ and can be thought of as candidates for characters of rational CFTs. Motivated by the observation that the same formulas appear as the half-indices of certain 3d ${\cal N}=2$ supersymmetric gauge theories, we perform a general search over low-rank 3d ${\cal N}=2$ abelian Chern-Simons matter theories which either flow to unitary TFTs or ${\cal N}=4$ rank-zero SCFTs in the infrared. These are exceptional classes of 3d theories, which are expected to support rational and $C_{2}$-cofinite chiral algebras on their boundary. We compare and contrast our results with Zagier’s and comment on a possible generalization of Nahm’s conjecture.
@@ 1 Introduction
Among various strategies, we focus on a series of works Nahm:1992sx ; Berkovich:1994es ; Nahm:1994vas ; Zagier:2007knq ; Nahm:2004ch , which highlighted the fact that a large class of 2d VOA characters can be represented as a Nahm sum $\mathbb{N}$ is a set of non-negative integers including 0..
@@
A natural question is the following: for which matrices $A$ do there exist $B$ and $C$ such that (3) becomes a modular function? W. Nahm attempted to answer this question by developing a conjecture which relates modular functions and torsion elements in the Bloch group. Utilizing this conjecture, Nahm, and separately D. Zagier, found numerous examples of triplets $(A,B,C)$ for which (3) is modular Nahm:2004ch ; Zagier:2007knq .
@@
where each of the $U(1)$ factors has precisely one charged chiral multiplet of charge 1222See recent works Jockers:2021omw ; Chung:2023qth ; Gang:2023rei ; Gang:2023ggt ; Ferrari:2023fez ; Creutzig:2024ljv ; Gaiotto:2024ioj , which focuses on a similar observation.. In this context, the matrix $A$ is identified with the mixed Chern-Simons level matrix $K$. As discussed extensively in recent literature Costello:2018fnz ; Costello:2018swh ; Costello:2020ndc ; Creutzig:2021ext ; Coman:2023xcq , the boundary vertex operator algebras of 3d supersymmetric gauge theories are generically non-rational. However, there are two important classes of exceptions:
@@ 2 Nahm sum formula and modular functions
which is obtained by examining the asymptotic behavior of the Nahm sum formula (3) as $q\rightarrow 1$. Nahm’s conjecture can be stated in a way that relates the solutions of (6) to the modularity of (3). This is achieved via two special functions known as the Rogers dilogarithm and the Bloch-Wigner function. The Rogers dilogarithm is defined on $10<x<$ 1 as
@@
This function is continuous for all $z\in\mathbb{C}$ and satisfies $D(z)=0$ for $z\in\mathbb{R}$. For $A$ positive definite and symmetric with entries in $\mathbb{Q}$, (6) has exactly one solution with all real algebraic entries between 0 and 1. We denote the solutions to (6) as $X^{(a)}_{i}$ with $a$ indexing the solutions and beginning at $X^{(0)}_{i}$, which is defined to be the special solution with $10<X^{(0)}_{i}<$ 1. With this information Nahm’s conjecture can be stated as follows [^3] This is not the original form of Nahm’s conjecture as discussed in Nahm:2004ch ; Zagier:2007knq for which a counterexample was found in vlasenko:2011 . The conjecture given here is equivalent to a weakened which is discussed in Calegari:2017itq ..
@@
The previous work on Nahm’s conjecture is best described as a bellwether rather than a precise target we would wish to replicate. For one, because we study Chern-Simons levels, all of our results are integer-valued matrices; a restriction that does not exist when approaching this problem from a purely mathematical point of view. On the other hand, when comparing our findings with those of Nahm and Zagier, we should expect agreement only when their $(A,B)$ are integral. [^5] The integrality of $B$ arises from the physical interpretation of the linear term in the half index. Further, the possible appearance of an additional phase in (11) indicates that we are working with a slight generalization of Nahm’s conjecture, where $L(x)$ is modified by a linear term in $\log(x)$ (See Sec. A.4 for a discussion). Therefore, we should expect to discover new matrices beyond those originating from Nahm’s conjecture. Indeed, this is what we find: any level matrix $K$ listed in Section 4 that is not also given in Table 1 either does not satisfy $(i)$ (and thus was ignored by Zagier) or does not satisfy the conditions of Nahm’s conjecture due to the more general form of (10). For these novel cases in Section 4, we find that there are negative contributions in the character summation formula, i.e. (10) does not reduce to (3) and thus does not contradict the partial proof of Nahm’s conjecture in Calegari:2017itq .
@@ 3.1 ${\cal N}=2$ abelian Chern-Simons matter theories
We consider a class of ${\cal N}=2$ abelian Chern-Simons matter theories, which are labeled by the Chern-Simons level matrix $K$ and a set of chiral operators $\{{\cal O}_{I}\}$
@@
Here ${\cal T}_{\Delta}$ is a free theory of a single chiral field $\Phi$ with background Chern-Simons level $-\frac{1}{2}$ Dimofte:2011ju and a Lagrangian density given by
@@
Here $\Sigma_{V}$ is the linear superfield containing the field strength of the background vector multiplet $V$ coupled to the $U(1)$ flavor symmetry. The theory also has $U(1)$ R-symmetry. The mixed CS levels between the R-symmetry and the flavor symmetry are summarized in Table 2.
@@
The notation $/U(1)^{r}_{K}$ denotes the supersymmetric gauging of the $U(1)^{r}$ flavor symmetry in the $({\cal T}_{\Delta})^{r}$ theory ($r$ copies of ${\cal T}_{\Delta}$) with mixed Chern-Simons level $K$. The gauge charge $Q_{ij}$ of the $i$-th chiral field $\Phi_{i}$ under the $j$-th $U(1)$ gauge symmetry is chosen to be $\delta_{ij}$. If we denote $\{{\cal V}_{i}\}$ to be the background vector multiplets for $(U(1)_{\text{Top.}})^{r}$ topological symmetry, then the Lagrangian density of the $({\cal T}_{\Delta})^{r}/U(1)^{r}_{K}$ theory is
@@
Denote $T_{1},\ldots,T_{r}$ as the charges of the $U(1)_{\text{top}}^{r}$ topological symmetry which exists in the absence of the superpotential deformation. The theory also has a $U(1)_{R}$ symmetry which can be mixed with $U(1)_{\text{top}}^{r}$. Let $R_{\vec{\mu}}$ with $\vec{\mu}\in\mathbb{R}^{r}$ be the R-charge at general mixing defined as
@@
In order to identify a candidate theory ${\cal T}[K,\{{\cal O}_{I}\}]$ which flows to an IR ${\cal N}=4$ rank-0 SCFT, we will require that there are $(r-1)$ 1/2-BPS CPOs in the superpotential ${\cal W}$, i.e. $N_{\cal O}=r-1$. This is because an ${\cal N}=4$ rank-0 theory should possess only one (non-R) $U(1)$ flavor symmetry when described in terms of an ${\cal N}=2$ subalgebra.777 However, it is possible for some ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ flow to an ${\cal N}=4$ rank-0 SCFT, with some UV symmetries decoupling in the IR. We will discuss such exceptional cases in section 4.3. In this case, the affine subspace is one-dimensional and can be represented as
$\mathfrak{M}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]=\{\vec{\mu}=\vec{\mu}_{0}+\nu\vec{a}\;:\;\vec{\nu}\in\mathbb{R}\}.$ (21)
The $\vec{\mu}_{0}$ is chosen such that $R_{\vec{\mu}_{0}}$ is the superconformal R-charge of the theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ which is determined by F-maximization as explained below. The vector $\vec{a}$ is orthogonal to the monopole fluxes $\mathbf{m}_{I}$ of ${\cal O}_{I}$ and its overall normalization is fixed by requiring that it is a primitive element in $\mathbb{Z}^{r}$ with first non-zero component positive. The theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ has $U(1)_{A}$ flavor symmetry with charge $A$ given by
$A=\vec{a}\cdot\vec{T}=\sum_{i=1}^{r}a_{i}T_{i}\;.$ (22)
If SUSY enhancement occurs, the $U(1)_{A}\times U(1)_{R}$ is expected to enhance to $SO(4)_{R}=SU(2)^{H}\times SU(2)^{C}$ R-symmetry. We identify
@@
and define the embedding as
@@
Here $J_{3}^{H}/J_{3}^{C}$ are the Cartan generators of $SU(2)^{H}/SU(2)^{C}$ normalized as $J_{3}\in\frac{1}{2}\mathbb{Z}$.
@@ 3.2 Conditions on supersymmetric partition functions
The main supersymmetric observable we compute is the superconformal index. For the theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$, the index is defined by
$\begin{split}{\cal I}_{\rm sci}(q,\eta,\nu)&=~{}\textrm{Tr}_{{\cal H}_{\rm rad}(S^{2})}(-1)^{R_{\nu}}q^{\frac{{R}_{\nu}}{2}+j_{3}}\eta^{A}.\end{split}$ (25)
Here, ${\cal H}_{\rm rad}(S^{2})$ denotes the radially quantized Hilbert-space and the index counts local operators (with signs). It can be computed by the integral formula Kim:2009wb ; Imamura:2011su
@@
The IR superconformal R-charge $\vec{\mu}_{0}$ of ${\cal T}[K,\{{\cal O}_{I}\}]$ is determined by the fact that
@@
is maximized at $\vec{\mu}=\vec{\mu}_{0}$ in the space of possible mixings $\mathfrak{M}[K,\{{\cal O}_{I}\}]$ Jafferis:2010un ; Jafferis:2011zi ; Closset:2012vg . ${\cal Z}_{S^{3}_{b}}^{K}$ is the squashed three-sphere partition function (with squashing parameter $b$) which can be written as the integral Hama:2011ea :
@@
Given a theory ${\cal T}[K,\{{\cal O}_{I}\}]$, we can compute each of the observables listed above and determine if the theory meets the following criteria:
@@
    There are $(r-1)$ linearly independent [^8] The set $\{(\mathbf{n}_{I},\mathbf{m}_{I})\in\mathbb{Z}^{2r}\}_{I=1}^{r-1}$ is linearly independent. 1/2 BPS CPOs $\{{\cal O}_{I}={\cal O}_{(\mathbf{n}_{I},\mathbf{m}_{I})}\}_{I=1}^{r-1}$ satisfying (17).
    Note that it is possible for ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ to flow to an $\mathcal{N}=4$ rank-0 SCFT due to excess $U(1)$ flavor symmetries decoupling in the IR. For simplicity, we focus on the case with $N_{\cal O}=r-1$.
@@
    If SUSY enhancement occurs, the ${\cal N}=2$ superconformal R-charge $R_{\nu=0}=R_{\vec{\mu}_{0}}=R_{*}+\vec{\mu}_{0}\cdot\vec{T}$ can only take values in $\frac{1}{2}\mathbb{Z}$ since $R_{\nu=0}=J_{3}^{C}+J_{3}^{H}$. Taking into account the fact that $R_{*}$ and $T_{i}$ are integer-valued, we arrive at the condition above. This condition is restrictive as $\vec{\mu}_{0}$ is determined by extremizing $|{\cal Z}^{K}_{S^{3}_{b=1}}(\vec{\mu})|$, which is in general a transcendental function of $\vec{\mu}$.
@@
    If ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ flows to an ${\cal N}=4$ SCFT, ${\cal I}_{\rm sci}(q,\nu=\pm 1,\eta=1)$ only receives contributions from Coulomb/Higgs branch operators and their descendants Razamat:2014pta . Thus the index must be trivial in this limit if the IR SCFT is rank-0. Condition $(ii)$ ensures that the IR theory is a non-trivial SCFT with no mass gap. We examine the gapped cases in Section 4.2.
When satisfied these conditions provide strong evidence that the theory flows to an ${\cal N}=4$ rank-zero SCFT.
In Section 4, we list the theories ${\cal T}[K,\{{\cal O}_{I}\}]$ which have a superconformal index satisfying conditions (a), (b), and (c). It is also possible to compute other supersymmetric observables, such as twisted supersymmetric partition functions on Seifert manifolds, to provide further consistency conditions. As discussed in detail in Appendix A.3, one can extract the modular data of the boundary VOA from the partition functions and this must satisfy a number of non-trivial relations. In particular, the fibering operator ~$\widetilde{{\cal F}}(u,v)$ evaluated at the Bethe vacua $u=\hat{u}$ are identified with the diagonal components of the modular $T$ matrices (see (LABEL:ST-HF_relations) and (LABEL:tH_and_tF)), which leads to the condition
~$\widetilde{{\cal F}}(\hat{u},v)^{N}=1\ ,$ (30)
for some integer $N$ and for all vacua $u=\hat{u}$. This gives a direct connection to Nahm’s conjecture as summarized in Section 2.

### v2

# 3d Topological Field Theories and Nahm Sum Formulas

@@ Abstract
It is known that a large class of characters of 2d conformal field theories (CFTs) can be written in the form of a Nahm sum. In Zagier:2007knq , D. Zagier identified a list of Nahm sum expressions that are modular functions under a congruence subgroup of $SL(2,\mathbb{Z})$, which can be thought of as candidates for characters of rational CFTs. Motivated by the observation that the same formulas appear as the half-indices of certain 3d ${\cal N}=2$ supersymmetric gauge theories, we perform a general search over low-rank 3d ${\cal N}=2$ abelian Chern-Simons matter theories which either flow to unitary TFTs or ${\cal N}=4$ rank-zero SCFTs in the infrared. These are exceptional classes of 3d theories, which are expected to support rational and $C_{2}$-cofinite chiral algebras on their boundary. We compare and contrast our results with Zagier’s and comment on a possible generalization of Nahm’s conjecture.
@@ 1 Introduction
[…]
@@
A natural question is the following: for which matrices $A$ do there exist $B$ and $C$ such that (3) becomes a modular function? W. Nahm attempted to answer this question by developing a conjecture which relates modular functions and torsion elements in the Bloch group. Utilizing this conjecture, Nahm and separately D. Zagier found numerous examples of triplets $(A,B,C)$ for which (3) is modular Nahm:2004ch ; Zagier:2007knq .
@@
where each of the $U(1)$ factors has precisely one charged chiral multiplet of charge 1.  [^2] See recent works Kucharski:2017ogk ; Ekholm:2020lqy ; Jockers:2021omw ; Chung:2023qth ; Gang:2023rei ; Gang:2023ggt ; Ferrari:2023fez ; Creutzig:2024ljv ; Gaiotto:2024ioj , which focuses on a similar observation. In this context, the $r\times r$ matrix $A$ is identified with the mixed Chern-Simons level matrix $K$. As discussed extensively in recent literature Costello:2018fnz ; Costello:2018swh ; Costello:2020ndc ; Creutzig:2021ext ; Coman:2023xcq , the boundary vertex operator algebras of 3d supersymmetric gauge theories are generically non-rational. However, there are two important classes of exceptions:
@@ 2 Nahm sum formula and modular functions
which is obtained by examining the asymptotic behavior of the Nahm sum formula (3) as $q\rightarrow 1$. Nahm’s conjecture can be stated in a way that relates the solutions of (6) to the modularity of (3). This is achieved via two special functions known as the Rogers dilogarithm and the Bloch-Wigner function. The Rogers dilogarithm is defined on $0<x<1$ as
@@
This function is continuous for all $z\in\mathbb{C}$ and satisfies $D(z)=0$ for $z\in\mathbb{R}$. For $A$ positive definite and symmetric with entries in $\mathbb{Q}$, (6) has exactly one solution with all real algebraic entries between 0 and 1. We denote the solutions to (6) as $X^{(a)}_{i}$ with $a$ indexing the solutions and beginning at $X^{(0)}_{i}$, which is defined to be the special solution with $0<X^{(0)}_{i}<1$. With this information Nahm’s conjecture can be stated as follows [^3] This is not the original form of Nahm’s conjecture as discussed in Nahm:2004ch ; Zagier:2007knq for which a counterexample was found in vlasenko:2011 . The conjecture given here is equivalent to a weakened which is discussed in Calegari:2017itq ..
@@
The previous work on Nahm’s conjecture is best described as a bellwether rather than a precise target we would wish to replicate. For one, because we study Chern-Simons levels, all of our results are integer-valued matrices; a restriction that does not exist when approaching this problem from a purely mathematical point of view. On the other hand, when comparing our findings with those of Nahm and Zagier, we should expect agreement only when their $(A,B)$ are integral [^5] The integrality of $B$ arises from the physical interpretation of the linear term in the half index.. Further, the possible appearance of an additional phase in (11) indicates that we are working with a slight generalization of Nahm’s conjecture, where $L(x)$ is modified by a linear term in $\log(x)$ (See Sec. A.4 for a discussion). Therefore, we should expect to discover new matrices beyond those originating from Nahm’s conjecture. Indeed, this is what we find: any level matrix $K$ listed in Section 4 that is not also given in Table 1 either does not satisfy $(i)$ (and thus was ignored by Zagier) or does not satisfy the conditions of Nahm’s conjecture due to the more general form of (10). For these novel cases in Section 4, we find that there are negative contributions in the character summation formula, i.e. (10) does not reduce to (3) and thus does not contradict the partial proof of Nahm’s conjecture in Calegari:2017itq .
@@ 3.1 ${\cal N}=2$ abelian Chern-Simons matter theories
We consider a class of ${\cal N}=2$ abelian Chern-Simons matter theory, which are labeled by the Chern-Simons level matrix $K$ and a set of chiral operators $\{{\cal O}_{I}\}$:
@@
Here ${\cal T}_{\Delta}$ is a free theory of a single chiral field $\Phi=(\phi,\psi)$ with background Chern-Simons level $\frac{1}{2}$ Dimofte:2011ju with a Lagrangian density given by
@@
Here $\Sigma_{V}$ is the linear superfield containing field strength of the background vector multiplet $V=(A_{\mu},\sigma,\lambda,D)$ coupled to the $U(1)$ flavor symmetry. The theory also has $U(1)$ R-symmetry. The mixed CS levels between the R-symmetry and the flavor symmetry are summarized in Table 2.
@@
The notation $U(1)^{r}_{K}$ in (13) denotes the supersymmetric gauging of the $U(1)^{r}$ flavor symmetry in the $({\cal T}_{\Delta})^{r}$ theory ($r$ copies of ${\cal T}_{\Delta}$) with mixed Chern-Simons level $K$, which is an $r\times r$ symmetric integer-valued matrix. The gauge charge $Q_{ij}$ of the $i$-th chiral field $\Phi_{i}$ under the $j$-th $U(1)$ gauge symmetry is chosen to be $\delta_{ij}$. If we denote $\{{\cal V}_{i}\}$ to be the background vector multiplets for $(U(1)_{\text{Top.}})^{r}$ topological symmetry, then the Lagrangian density of the $({\cal T}_{\Delta})^{r}/U(1)^{r}_{K}$ theory is
@@
Accounting for the background Chern–Simons level $k_{FF}=-\frac{1}{2}$ in the $\mathcal{T}_{\Delta}$ theory (2), the UV effective Chern–Simons level for the $U(1)^{r}$ vector multiplet is
$K-\frac{1}{2}I\ .$ (16)
@@
Let us denote $T_{1},\ldots,T_{r}$ by the charges under the $U(1)_{\text{top}}^{r}$ topological symmetry which exists in the absence of the superpotential deformation. The theory also has a $U(1)_{R}$ symmetry which can be mixed with $U(1)_{\text{top}}^{r}$. Let $R_{\vec{\mu}}$ with $\vec{\mu}\in\mathbb{R}^{r}$ be the R-charge at general mixing
@@
In order to identify a candidate theory ${\cal T}[K,\{{\cal O}_{I}\}]$ which flows to an IR ${\cal N}=4$ rank-0 SCFT, we will require that there are (r1)(r-1) 1/2-BPS CPOs in the superpotential $\cal W$}, i.e. $N_{\cal O}=r-1$. This is because an ${\cal N}=4$ rank-0 theory should possess only one (non-R) $U(1)$ flavor symmetry when described in terms of an ${\cal N}=2$ subalgebra.777 However, it is possible for some ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ flow to an ${\cal N}=4$ rank-0 SCFT, with some UV symmetries decoupling in the IR. We will discuss such exceptional cases in section 4.3. In this case, the affine subspace is one-dimensional and can be represented as
$\mathfrak{M}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]=\{\vec{\mu}=\vec{\mu}_{0}+\nu\vec{a}\;:\;\nu\in\mathbb{R}\}.$ (22)
The $\vec{\mu}_{0}$ is chosen such that $R_{\vec{\mu}_{0}}$ is the superconformal R-charge of the theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ which is determined by F-maximization as explained below. The vector $\vec{a}$ is orthogonal to the monopole fluxes $\mathbf{m}_{I}$ of ${\cal O}_{I}$ and its overall normalization is fixed by requiring that it is a primitive element in $\mathbb{Z}^{r}$ with first non-zero component positive. The theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ has a $U(1)_{A}$ flavor symmetry with charge $A$ given by
$A=\vec{a}\cdot\vec{T}=\sum_{i=1}^{r}a_{i}T_{i}\;\ ,$ (23)
which will be later identified with the part of the R-symmetry in the IR ${\cal N}=4$ algebra.
Topological twisting
If supersymmetry enhances at the superconformal fixed point, the $U(1)_{A}\times U(1)_{R}$ symmetry extends to the full $SO(4)_{R}=SU(2)^{H}\times SU(2)^{C}$ R-symmetry of the ${\cal N}=4$ algebra. This allows the topological A/B-twist, defined by the twisting homomorphism
$SU(2)^{E}\rightarrow\text{diag}(SU(2)^{E}\times SU(2)^{H})\ ,\quad SU(2)^{E}\rightarrow\text{diag}(SU(2)^{E}\times SU(2)^{C})\ ,$ (24)
for the $A$- and $B$-twist, respectively. Each twist preserves a real scalar supercharge, denoted $Q_{A}$ and $Q_{B}$. The local operators of the superconformal theory that survive in the cohomology of $Q_{A}$ and $Q_{B}$ correspond to Coulomb branch operators and Higgs branch operators, respectively.
From the point of view of the ${\cal N}=2$ superconformal algebra, there exists a distinguished $U(1)$ flavor symmetry called the axial symmetry
@@
where $J_{3}^{H}/J_{3}^{C}$ are the Cartan generators of $SU(2)^{H}/SU(2)^{C}$ R-symmetry of ${\cal N}=4$ algebra, normalized so that $J_{3}\in\frac{1}{2}\mathbb{Z}$. We then define a one-parameter deformation of the $U(1)$ superconformal R-symmetry by mixing with the axial symmetry
@@
The two special values of $\nu$ play an important role:
$R_{\nu=1}=2J_{3}^{C}\ ,\qquad R_{\nu=-1}=2J_{3}^{H}\ ,$ (27)
as they coincide with the Cartan generators of the R-symmetry that defines the topological $B$- and $A$-twist respectively.
@@ 3.2 Conditions on supersymmetric partition functions
The main observable we consider in this paper is the superconformal index. For the theory ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$, we define
$\begin{split}{\cal I}_{\rm sci}(q,\eta,\nu)&=~\textrm{Tr}_{{\cal H}(S^{2})}(-1)^{R_{\nu}}q^{\frac{{R}_{\nu}}{2}+j_{3}}\eta^{A}\ ,\end{split}$ (28)
which counts local operators graded by their twisted spins
$j_{\text{twisted}}=\frac{1}{2}R_{\nu}+j_{3}\ .$ (29)
As discussed around eq. (26), the parameter $\nu$ is defined so that $R_{\nu=0}$ corresponds to the superconformal R-symmetry at the fixed point, determined by F-maximization. Specializing to $\nu=+1/\nu=-1$ and $\eta=1$, the index computes the dimension of the space of Higgs/Coulomb branch local operators, graded by the twisted spins $J_{3}^{C,H}+j_{3}$,
${\cal I}_{\rm sci}(q,1,1)=~\textrm{Tr}_{{\cal H}(S^{2})}(-1)^{2J_{3}^{C}}q^{J_{3}^{C}+j_{3}}\ ,\quad{\cal I}_{\rm sci}(q,1,-1)=~\textrm{Tr}_{{\cal H}(S^{2})}(-1)^{2J_{3}^{H}}q^{J_{3}^{H}+j_{3}}\ ,$ (30)
which are also known as the Higgs/Coulomb branch Hilbert series. These quantities coincide with the $S^{2}\times S^{1}$ partition function of the B- and A-twisted theory, counting the local operators in each TFT.
The index can be computed by the integral formula Kim:2009wb ; Imamura:2011su
@@
The IR superconformal R-charge $\vec{\mu}_{0}$ of $T[K,\{{\cal O}_{I}\}]$ is determined by the fact that
@@
is maximized at $\vec{\mu}=\vec{\mu}_{0}$ in the space of possible mixings, $\mathfrak{M}[K,\{{\cal O}_{I}\}]$ Jafferis:2010un ; Jafferis:2011zi ; Closset:2012vg . ${\cal Z}_{S^{3}_{b}}^{K}$ is the squashed three-sphere partition function (with squashing parameter $b$) which can be written as the integral Hama:2011ea :
@@
Given a theory ${\cal T}[K,\{{\cal O}_{I}\}]$, we can compute each of the observables listed above and determine if the theory meets the following criteria. When satisfied these conditions provide strong evidence that the theory flows to an ${\cal N}=4$ rank-zero SCFT.
@@
    There are (r1)(r-1) linearly independent [^8] The set $\{(\mathbf{n}_{I},\mathbf{m}_{I})\in\mathbb{Z}^{2r}\}_{I=1}^{r-1}$ is linearly independent. 1/2 BPS CPOs $\{{\cal O}_{I}={\cal O}_{(\mathbf{n}_{I},\mathbf{m}_{I})}\}_{I=1}^{r-1}$ satisfying (18).
    Note that, as mentioned previously, it is possible for ${\cal T}[K,\{{\cal O}_{I}\}_{I=1}^{N_{\cal O}}]$ with $N_{\cal O}<r-1$ to flow to an $\mathcal{N}=4$ rank-0 SCFT due to excess $U(1)$ flavor symmetries decoupling in the IR. For simplicity, we focus on the case with $N_{\cal O}=r-1$.
@@
    If SUSY enhancement occurs, the ${\cal N}=2$ superconformal R-charge $R_{\nu=0}=R_{\vec{\mu}_{0}}=R_{*}+\vec{\mu}_{0}\cdot\vec{T}$ can only take values in $\frac{1}{2}\mathbb{Z}$ since $R_{\nu=0}=J_{3}^{C}+J_{3}^{H}$. Taking into account the fact that $R_{*}$ and $T_{i}$ are integer-valued, we arrive at the condition above. This condition is quite restrictive as $\vec{\mu}_{0}$ is determined by extremizing $|Z^{K}_{S^{3}_{b=1}}(\vec{\mu})|$, which is in general a transcendental function of $\vec{\mu}$.
@@
    If $T[K,\{{\cal O}_{I}\}_{I=1}^{r-1}]$ flows to an ${\cal N}=4$ SCFT, ${\cal I}_{\rm sci}(q,\nu=+1/\nu=-1,\eta=1)$ only receives contributions from Higgs/Coulomb branch operators and their descendants Razamat:2014pta . Thus the index must be trivial in this limit if the IR SCFT is rank-0. Condition $(ii)$ ensures that the IR theory is a non-trivial SCFT with no mass gap. We examine the gapped cases in Section 4.2.
In Section 4, we list the theories $T[K,\{{\cal O}_{I}\}]$ which have a superconformal index satisfying conditions (a), (b), and (c). It is also possible to compute other supersymmetric observables, such as twisted supersymmetric partition functions on Seifert manifolds, to provide further consistency conditions. As discussed in detail in Appendix A.3, one can extract the modular data of the boundary VOA from the partition functions and this must satisfy a number of non-trivial relations. In particular, the fibering operator ${\cal F}(u,v)$ evaluated at the Bethe vacua $u=\hat{u}$ are identified with the diagonal components of the modular $T$ matrices, see (LABEL:ST-HF_relations) and (LABEL:tH_and_tF), which leads to the condition
${\cal F}(\hat{u},v)^{N}=1\ ,$ (35)
for some integer $N$ and for all the vacua $u=\hat{u}$. This gives a direct connection to Nahm’s conjecture as summarized in Section 2.